# Psycholinguistic Analyses in Software Engineering Text: A Systematic Literature Review

### v1

@@ Abstract
Conclusion: We highlight the potential of psycholinguistic tools and their limitations, and present new use cases for advancing the research of human factors in SE (e.g., bias in human-LLM conversations).
@@ 1 Introduction
Understanding human factors in software engineering (SE) is critical to improving collaboration, decision making, and productivity in development teams. The vast amount of textual data generated by stakeholders, developers, and users – ranging from code reviews to chats – contains valuable insight into the psychological and emotional states of those involved in software projects. Understanding these factors can reveal key issues in software teams, such as job satisfaction, project participation, and developer attrition [1, 2, 3]. However, leveraging the full potential of textual data requires the use of specialized tools capable of accurately analyzing unique language patterns and vocabulary used in software engineering [4, 5].
Understanding human factors in software engineering (SE) is critical to improving collaboration, decision making, and productivity in development teams. The vast amount of textual data generated by stakeholders, developers, and users – ranging from code reviews to chats – contains valuable insight into the psychological and emotional states of those involved in software projects. Understanding these factors can reveal key issues in software teams, such as job satisfaction, project participation, and developer attrition [1, 2, 3]. However, leveraging the full potential of textual data requires the use of specialized tools capable of accurately analyzing unique language patterns and vocabulary used in software engineering [4, 5].
@@
    We identified 17 studies that conducted formal evaluations of LIWC in SE contexts. The evaluation methods varied, including comparisons with ground-truth data and reliable psychometric tools from psychology. These evaluations varied in scope but provided validation of LIWC’s effectiveness in SE research.
@@ 2.3 Linguistic Inquiry and Word Count (LIWC)
James W. Pennebaker et al. initially developed Linguistic Inquiry and Word Count (LIWC) to automate parts of the text analysis process they intended to use for analyzing essays written by people who have experienced trauma. Since then, LIWC has been used across different disciplines to analyze text and, among other use-cases, extract psychological information from the text [22].
In a top-down fashion, LIWC’s dictionaries categorize words into predetermined categories, mapping them to important psychological constructs and theories. The categories cover a wide range of linguistic and psychological dimensions, including cognition, affect, social processes, culture, and motives, to name a few. Table LABEL:tab:all_categories lists all the categories directly mentioned in our selected studies that are still available in the 2022 version of the LIWC. Along the categories, we include the explanations or common words listed in each category.
Table 1: LIWC-22 Categories Mentioned in our Subjects of Study
@@
Word count Total word count
Analytical thinking Perceived honesty, genuineness
Clout Language of leadership, status
Emotional tone Degree of positive (negative) tone
Big words Percent words 7 letters or longer
Total function words the, to, and, I
1st person plural we, our, us, lets
@@
Impersonal pronouns that, it, this, what
Numbers one, two, first, once
Auxiliary verbs is, was, be, have
Conjunctions and, but, so, as
Common verbs is, was, be, have
Achievement work, better, best, working
Cognitive processes but, not, if, or, know
Leisure game*, fun, play, party*
Money business*, pay*, price*, market*
Past focus was, had, were, been
Present focus is, are, I’m, can
Future focus will, going to, have to, may
Time when, now, then, day
Visual see, look, eye*, saw
Auditory sound*, heard, hear, music
Fillers rr*, wow, sooo*, youknow
Authentic Perceived honesty, genuineness
Words per sentence Average words per sentence
Dictionary words Percent words captured by LIWC
1st person singular I, me, my, myself
2nd person you, your, u, yourself
@@
Based on the word count, i.e., the frequency of specific words in a body of text, and the percentages of words falling into each of the categories, LIWC provides insight into psychological and emotional content in text. Over the years, various studies have validated the correlation between LIWC’s output and a wide range of psychological concepts. In a 2010 publication, Tausczik & Pennebaker pointed out these correlations by referencing studies that used tools such as surveys, personality assessments, and self-reports to analyze participants’ linguistic patterns [22]. For instance, they highlight how the use of words under the 1st person singular category e.g., “I”, “me”, and “mine” reflects self-focus. Psychological concepts like self-reflection, insecurity, anxiety, and even depression have been shown to correlate with increased use of 1st person singular words in various contexts, as individuals experiencing these states tend to center their attention inwardly, which is reflected in their language use. These findings were established through empirical research that connected participants’ word usage with their self-reported psychological states, providing strong evidence of the links between language patterns and psychological phenomena.
The value for most LIWC categories are calculated based on the percentage of the words in the text that fall under them. However, LIWC’s summary variables, or the four categories of analytical thinking, clout, authenticity, and emotional tone are, in particular, derived from the percentage of words in other LIWC categories. These summary variables stand as the only non-transparent dimensions within the LIWC program, as the precise calculations leading to the values assigned to these categories remain unknown.
Since the initial release of LIWC, multiple versions of it have been created, each characterized by a larger dictionary and a more extensive set of interpretable dimensions. LIWC2015, for instance uses dictionaries with as many as 4500 words [23], while LIWC-22’s internal dictionary contains 12000 words [24]. Notably, new functionalities are added to the 2022 version of LIWC that have been absent in the previous versions. For instance, LIWC-22 offers narrative arc, word frequency and word clouds, and a dictionary workbench designed to streamline the cumbersome process behind the creation of custom dictionaries [24]. Lists of the dimensions and capabilities included in each version of LIWC can be found in their corresponding published journal articles or manuals, e.g., https://www.liwc.app/help/psychometrics-manuals.
@@ 3.1 Systematic Literature Reviews of Behavioral Analyses in Software Engineering
Several systematic literature reviews were conducted to identify the use of behavioral analyses in SE. The scope of studies incorporating behavioral analysis in SE is extensive, covering topics such as sentiment and emotion detection [13], personality assessment [25], and job satisfaction [3]. Researchers employ a wide range of methodologies, from traditional tools like questionnaires and interviews to automated Machine Learning (ML) models, to answer their research questions. This diversity highlights the vast potential for further exploration to deepen our understanding of the human element in the software engineering process.
@@
Numerous studies have also explored the impact of various human factors within development environments, along with methods for their detection. Lin et al. conducted a systematic literature review focusing on opinion mining in software development, including tasks such as sentiment analysis, emotion detection, and politeness detection [13]. Another survey study looked at empirical research on developer emotions between 2005 and 2018 [27]. Felipe et al. conducted a similar review, examining the psychometric tools used to infer personality types in software engineering research [25]. Studies of these nature frequently identify LIWC as a tool for detecting the psychological concepts and human factors they examine. However, no study has yet comprehensively examined the role of LIWC across the entire body of software engineering research.
@@ 3.2 Use of LIWC Outside Software Engineering
The work of Pennebaker et al. reflects LIWC’s versatile use-cases [28]. For instance, LIWC has been frequently used to assess the emotional content of text, performing tasks such as sentiment analysis and the detection of specific emotions [29]. Additionally, it has been used to assess the personality traits of individuals who contribute to the text [25]. In psychology research, many have used LIWC to detect depression [6, 30] and to study how the language use of couples reltes to their marital satisfaction and the overall heath of the marriage [8, 9]. LIWC has also been used to analyze text with political content to assess the public’s opinion on political topics and to find common linguistic patterns in speech of famous politicians [31, 32, 33]. Other researchers have analyzed psychological concepts within a wide range of texts, including suicidal poems [34], Al-Qaeda transcripts [35], and social media posts written by users dealing with the impacts of the Covid-19 pandemic [36]. The variety and quantity of these studies have showcased LIWC’s unparalleled ability to analyze human emotions, thoughts, and intentions encoded within language.
@@ 4 Review Methodology
The goal of this paper is to review and analyze the software engineering research that uses LIWC. Following the review study methodology by Kitchenham et al. [37], this section outlines our approach toward identifying relevant research papers. An overview of our approach for selecting the relevant papers is shown in Figure 1.
@@
To select the papers for our survey, we first curated a query to retrieve as many relevant research papers in digital libraries as possible. We examined the retrieved papers closely (first title and abstract and then the full text) to filter out any irrelevant or inappropriate studies based on a set of specific inclusion and exclusion criteria (detailed description in Table 3). This step was followed by forward and backward snowballing in order to locate any papers that our initial query may have missed. Finally, the newly collected papers were again manually filtered, leaving us with the final set of papers that we analyzed in this study (n=43).
@@ 4.1 Online Query
Through a process of trial and error, coupled with insights from prior literature regarding suitable phrases for the retrieval of the software engineering papers [13], we curated the following search query:
@@
The initial part of the query, (“liwc” OR “linguistic inquiry and word count”), is designed with the intent of inclusively capturing papers that reference the complete name of the targeted tool or its widely recognized abbreviation, “LIWC”. The latter component of our query, (“software engineering”), serves as a filter, enabling the identification of papers discussing subjects related to the field of software engineering.
@@
4.2 Paper Selection
To support the manual paper selection process, we first established a precise set of inclusion and exclusion criteria, as shown in Table 3. All the papers included in this study must have undergone peer review and have been published in conferences, workshops, or journals. These papers primarily focus on topics related to software development and use LIWC in their analyses.
@@ 4.2 Paper Selection
In total, our paper selection process led to the identification of 43 studies (36 in the first iteration + 7 through snowballing) that are the subjects of this survey. Table 4 illustrates the distribution of our chosen papers based on their publication years. The first software engineering paper using LIWC, a pioneering paper by Rigby et al., was published in 2007 [S27]. Since its publication, Rigby et al.’s work has inspired many software engineering studies that have used LIWC to directly analyze many facets of developers’ communications. After 2013, the number of software engineering publications that leveraged LIWC in their studies has noticeably grown. As of 2023, LIWC continues to feature prominently in software engineering studies, highlighting the community’s ongoing interest and the relevance of this tool in the research area.
@@ 5.1 RQ1: How was LIWC employed in SE studies, and for what specific activities and purposes?
Through answering RQ1, we also aim to better understand the types of LIWC applications within the Software Engineering (SE) research. We pin down the purposes LIWC has fulfilled in each study and, additionally, we find out whether it was used to directly interpret the text and study psychological concepts or if it was used indirectly as explained in the remainder of this section. In Table LABEL:tab:applications, we list all the direct and indirect uses of LIWC along with the categories used for each performed activity and task. Next, we describe both scenarios in detail.

### v2

@@ Abstract
Conclusion: We highlight the potential of psycholinguistic tools and their limitations, and present new use cases for advancing research on human factors in SE (e.g., bias in human-LLM conversations).
@@ 1 Introduction
Understanding human factors in SE is critical to improving collaboration, decision making, and productivity in development teams. The vast amount of textual data generated by stakeholders, developers, and users – ranging from code reviews to chats – contains valuable insight into the psychological and emotional states of those involved in software projects. Understanding these factors can reveal key issues in software teams, such as job satisfaction, project participation, and developer attrition [1, 2, 3]. However, leveraging the full potential of textual data requires the use of specialized tools capable of accurately analyzing unique language patterns and vocabulary used in software engineering [4, 5].
@@
    We identified 17 studies that conducted formal evaluations of LIWC in SE contexts. In our review, evaluation refers to any effort by researchers to assess LIWC’s effectiveness or reliability within their studies. These evaluations varied in scope but provided validation of LIWC’s role in SE research.
    We identified 17 studies that conducted evaluations of LIWC in SE contexts. In our review, evaluation refers to any effort by SE researchers to assess LIWC’s effectiveness or reliability. Methods included comparisons with ground-truth data and other forms of methodological validation. These evaluations varied in scope but provided evidence supporting LIWC’s use in SE research.
@@ 2.3 Linguistic Inquiry and Word Count (LIWC)
James W. Pennebaker et al. initially developed Linguistic Inquiry and Word Count (LIWC)  [^1] https://www.liwc.app to automate parts of the text analysis process they intended to use for analyzing essays written by people who have experienced trauma. Since then, LIWC has been used across different disciplines to analyze text and, among other use-cases, extract psychological information from the text [22].
In a top-down fashion, LIWC’s dictionaries categorize words into predetermined categories, mapping them to important psychological constructs and theories. The categories cover a wide range of linguistic and psychological dimensions, including cognition, affect, social processes, culture, and motives, to name a few. Table LABEL:tab:all_categories lists all the categories directly mentioned in our selected studies that are still available in the 2022 version of the LIWC.
To improve clarity, we have divided LIWC categories in Table LABEL:tab:all_categories into three subsections as follows: (1) Summary Variables, (2) Text Statistics, and (3) Lexical Variables. The values for Text Statistics and Lexical Variables are calculated based on the percentage of the words in the text that fall under them. However, Summary variables are derived from the percentage of words in other LIWC categories. The Summary Variables stand as the only non-transparent dimensions within the LIWC program, as the precise calculations leading to the values assigned to these categories remain unknown. In Table LABEL:tab:all_categories, we include the explanations or common words listed in each category. For example, under Lexical Variables section, “Drives” category includes words that are intended to capture motivational language related to affiliation, achievement, and cognitive engagement (e.g., we, know, work).
Based on the word count, i.e., the frequency of specific words in a body of text, and the percentages of words falling into each of the categories, LIWC provides insight into psychological and emotional content in text. Over the years, various studies have validated the correlation between LIWC’s output and a wide range of psychological concepts. In a 2010 publication, Tausczik & Pennebaker pointed out these correlations by referencing studies that used tools such as surveys, personality assessments, and self-reports to analyze participants’ linguistic patterns [22].
Table 1: LIWC-22 Categories Mentioned in Our Study Subjects.
@@
Summary Variables
Analytical thinking Measure of formal/logical thinking
Clout Measure of language indicating confidence and status
Authentic Measure of perceived honesty and genuineness
Emotional tone Measure of of positive vs. negative affect
Text Statistics
Word count Total number of words
Words per sentence Average words per sentence
Big words Percentage of words 7 letters or longer
Dictionary words Percentage of words matched in LIWC dictionary
Total function words Total number of function words (e.g., the, to, and, I)
Lexical Variables
1st person singular I, me, my, myself
1st person plural we, our, us, let’s
2nd person you, your, u, yourself
@@
Conjunctions and, but, so, as
Auxiliary verbs is, was, be, have
Common verbs is, was, be, have
Impersonal pronouns that, it, this, what
Numbers one, two, first, once
@@
Achievement work, better, best, working
Leisure game*, fun, play, party*
Money business*, pay*, price*, market*
Time when, now, then, day
Past focus was, had, were, been
Present focus is, are, I’m, can
Future focus will, going to, have to, may
Visual see, look, eye*, saw
Auditory sound*, heard, hear, music
Fillers rr*, wow, sooo*, youknow
@@
For instance, they highlight how the use of words under the 1st person singular category e.g., “I”, “me”, and “mine” reflects self-focus. Psychological concepts like self-reflection, insecurity, anxiety, and even depression have been shown to correlate with increased use of 1st person singular words in various contexts, as individuals experiencing these states tend to center their attention inwardly, which is reflected in their language use. These findings were established through empirical research that connected participants’ word usage with their self-reported psychological states, providing strong evidence of the links between language patterns and psychological phenomena.
Since the initial release of LIWC, multiple versions of it have been created, each characterized by a larger dictionary and a more extensive set of interpretable dimensions. LIWC2015, for instance, uses dictionaries with as many as 4500 words [23], while LIWC-22’s internal dictionary contains 12000 words [24]. Notably, new functionalities were added to the 2022 version of LIWC that have been absent in the previous versions. For instance, LIWC-22 offers narrative arc, word frequency and word clouds, and a dictionary workbench designed to streamline the cumbersome process behind the creation of custom dictionaries [24]. Lists of the dimensions and capabilities included in each version of LIWC can be found in the corresponding published journal articles or manuals, e.g., https://www.liwc.app/help/psychometrics-manuals.
@@ 3.1 Systematic Literature Reviews of Behavioral Analyses in Software Engineering
Several SLRs were conducted to identify the use of behavioral analyses in SE. The scope of studies incorporating behavioral analysis in SE is extensive, covering topics such as sentiment and emotion detection [13, 25, 26, 27, 28], personality assessment [29, 30, 31], and job satisfaction [3, 32]. Researchers employ a wide range of methodologies, from traditional tools like questionnaires and interviews to automated Machine Learning (ML) models, to answer their research questions. This diversity highlights the vast potential for further exploration to deepen our understanding of the human element in the software engineering process.
@@
Numerous studies have also explored the impact of various human factors within development environments, along with methods for their detection. Lin et al. conducted a systematic literature review (SLR) focusing on opinion mining in software development, including tasks such as sentiment analysis, emotion detection, and politeness detection [13]. Another survey study looked at empirical research on developer emotions between 2005 and 2018 [34]. Felipe et al. conducted a similar review, examining the psychometric tools used to infer personality types in software engineering research [29]. Studies of this nature frequently identify LIWC as a tool for detecting the psychological concepts and human factors they examine. To date, however, no prior work has systematically reviewed how LIWC itself has been adopted, evaluated, and applied across the entire body of software engineering research. While other SLRs focus on broader psychological or behavioral techniques, our study is the first to concentrate exclusively on the role and methodological implications of LIWC within software engineering.
@@ 3.2 Use of LIWC Outside Software Engineering
The work of Pennebaker et al. reflects LIWC’s versatile use-cases [35]. For instance, LIWC has been frequently used to assess the emotional content of text, performing tasks such as sentiment analysis and the detection of specific emotions [36]. Additionally, it has been used to assess the personality traits of individuals who contribute to the text [29]. In psychology research, many have used LIWC to detect depression [6, 37] and to study how the language use of couples relates to their marital satisfaction and the overall health of the marriage [8, 9]. LIWC has also been used to analyze text with political content to assess the public’s opinion on political topics and to find common linguistic patterns in speech of famous politicians [38, 39, 40]. Other researchers have analyzed psychological concepts within a wide range of texts, including suicidal poems [41], Al-Qaeda transcripts [42], and social media posts written by users dealing with the impacts of the Covid-19 pandemic [43]. The variety and quantity of these studies have showcased LIWC’s unparalleled ability to analyze human emotions, thoughts, and intentions encoded within language.
Despite LIWC’s wide adoption across many disciplines, and its frequent appearance in SE studies, no existing review has examined its specific applications, limitations, and potential, specifically within the context of software engineering. This gap motivates the need for a focused SLR that examines LIWC’s contributions to SE research, an effort we undertake in this study.
@@ 4 Review Methodology
The goal of this paper is to review and analyze the software engineering research that uses LIWC. Following the review study methodology by Kitchenham et al. [44], this section outlines our approach toward identifying relevant research papers. An overview of our approach for selecting the relevant papers is shown in Figure 1. Additionally, we address PRISMA items in Kitchenham et al.’s checklist for secondary studies to ensure alignment with the systematic review standards [45]. For further reference, please find our PRISMA 2020 checklist attached in the Appendix B.
@@
To select the papers for our survey, we first curated a query to retrieve as many relevant research papers in digital libraries as possible. We examined the retrieved papers closely (first title and abstract and then the full text) to filter out any irrelevant or inappropriate studies based on a set of specific inclusion and exclusion criteria (detailed description in Table 3). This step was followed by forward and backward snowballing in order to locate any papers that our initial query may have missed. Specifically, we adopted a hybrid approach that combined comprehensive keyword-based searches with both forward and backward snowballing, as supported by recent SLR methodologies [46, 13, 34]. Finally, the newly collected papers were again manually filtered, leaving us with the final set of papers that we analyzed in this study (n=43). The remainder of this section explains each step of this process in more detail.
@@ 4.1 Online Query
To ensure a comprehensive retrieval of relevant studies, we refined our search query iteratively through a process of trial and validation. Initially, we started with broad queries such as “liwc” OR ”linguistic inquiry and word count”, which retrieved a large number of papers across multiple digital libraries, but many were unrelated to software engineering. We then progressively refined the query by adding domain-specific terms, such as “software”, “software engineering”, and “developer” OR “development”, to filter studies focused on software engineering applications. For instance, adding “software engineering” significantly reduced irrelevant results while maintaining coverage of key studies. Additionally, we explored broader psycholinguistic and behavioral software engineering terms, such as “psycholinguistic” OR ”psychometric” AND ”text analysis” AND ”software engineering”, to identify studies that might not explicitly reference LIWC but were relevant to our research scope. The final query (“liwc” OR ”linguistic inquiry and word count”) AND (”software engineering”) retrieved the highest number of relevant papers possible. A full record of all queries, retrieval counts, and query refinements is included in our replication package for transparency and reproducibility.
This process, coupled with insights from prior literature regarding suitable phrases for the retrieval of the software engineering papers [13], led us to select the following search query:
@@
The initial part of the query, (“liwc” OR “linguistic inquiry and word count”), is designed with the intent of inclusively capturing papers that reference the complete name of the targeted tool or its widely recognized abbreviation, “LIWC”. The latter component of our query, (“software engineering”), serves as a filter, enabling the identification of papers discussing subjects related to the field of software engineering while not posing as much limitations as terms such as “developer”.
@@
4.2 Paper Selection and Eligibility
To support the manual paper selection process, we first established a precise set of inclusion and exclusion criteria in Table 3, following the PRISMA framework’s requirements [45]. All the papers included in this study must have undergone peer review and have been published in conferences, workshops, or journals. These papers primarily focus on topics related to software development and use LIWC in their analyses. Additionally, two authors contributed to this manual selection process. The first author used the inclusion and exclusion criteria to identify relevant papers. After each of the two review phases, the second author independently examined the final list, verifying whether each paper’s inclusion aligned with the criteria. Discrepancies or uncertainties were discussed between the two authors until full agreement was reached. This step added a layer of rigor and helped ensure a consistent application of our selection criteria.
@@ 4.2 Paper Selection and Eligibility
In total, our paper selection process led to the identification of 43 studies (36 in the first iteration + 7 through snowballing) that are the subjects of this survey. The complete set of final papers used throughout this work is listed in Appendix A, where references to each study are provided. Henceforth, we refer to these papers with “S” prefixes e.g., [S1], [S2], etc. Table 4 illustrates the distribution of our chosen papers based on their publication years. The first software engineering paper using LIWC was published in 2007 [S27]. Since then several software engineering studies have used LIWC to directly analyze many facets of developers’ communications. After 2013, the number of software engineering publications that leveraged LIWC in their studies has noticeably grown. As of 2023, LIWC continues to feature prominently in software engineering studies, highlighting the community’s ongoing interest and the relevance of this tool in the research area.
@@
4.3 Data Extraction and Coding
In this section, we describe the approach used to extract and code data from each paper included in our review. After retrieving the papers from the sources discussed above, we automatically collected basic metadata for each paper, including its title, author(s), publication year, and venue i.e., journal or conference. This information, along with the papers themselves, is made available in our replication package  [^2] https://figshare.com/articles/dataset/Psycholinguistic_Analyses_in_Software_Engineering_Text_A_Systematic_Literature_Review/28774319.
In line with SLR requirements and in order to systematically analyze the content of each paper, we created a data extraction form relevant to our research questions. We used this form to record data such as the LIWC categories utilized, the datasets analyzed, the specific software engineering activities addressed and the behavioral software engineering concepts investigated. The list of all data fields required is presented below:
  1. 1.
    Paper title
  2. 2.
    Authors
  3. 3.
    Year of publication
  4. 4.
    Publication venue (journal/conference)
  5. 5.
    Purpose for using LIWC
  6. 6.
    LIWC utilization: direct interpretation or indirect use
  7. 7.
    LIWC categories used
  8. 8.
    Dataset(s) analyzed
  9. 9.
    Software engineering activities studied
  10. 10.
    Behavioral software engineering concepts addressed
  11. 11.
    Evaluation methods used to assess LIWC’s effectiveness
  12. 12.
    Mentioned limitations or concerns with using LIWC
For identifying the software engineering activities and behavioral software engineering (BSE) concepts, we adopted a deductive coding approach [54], applying established taxonomies as discussed in Sections 5.2 and 5.3. For other data extraction tasks, such as identifying the purpose of using LIWC, types of data analyzed, and concerns reported by authors, we employed an inductive coding approach [54], developing categories iteratively based on recurring patterns in the reviewed papers. Two authors independently coded all 43 studies. Following Kitchenham et al.’s guidelines for systematic reviews [44], we calculated Cohen’s Kappa coefficients for each coding dimension to assess inter-rater reliability. The resulting Cohen’s Kappa coefficients were: 0.61 for both software engineering activities and BSE concepts, and 1.00 for determining the direct/indirect use-cases of LIWC, the purpose for using LIWC, datasets used, evaluation techniques, and concerns or limitations. Where initial agreement was below the threshold of substantial agreement (i.e., less than 0.6), the coding was refined through discussion until substantial agreement was achieved across all dimensions, following the guidelines proposed by Landis and Koch [55]. The slightly lower agreement for SE activities and BSE concepts can be attributed to the inherent complexity and occasional overlap of categories and labels within each of these dimensions’ taxonomies, which may lead to subtle differences in interpretation. Nonetheless, the final levels of agreement, each above 0.6, meet the standard for substantial inter-rater reliability.
@@ 5.1.1 Software Engineering Activities
5.1.2 Direct and Indirect Use of LIWC
As part of answering RQ1, we examined how LIWC has been used across SE studies and identified two overall usage patterns: direct and indirect. In Table LABEL:tab:applications, we list all the direct and indirect uses of LIWC along with the categories used for each performed activity and task. We elaborate on both usage types below.
Among all the included studies in our survey, 19 used LIWC to directly analyze text. As the name suggests, direct use of LIWC refers to the cases in which researchers used LIWC’s outputs directly to interpret and gain insights from the textual data without any further processing. These use-cases frequently encompass tasks such as detecting developer emotions (e.g., anger, negative emotion, and anxiety) [S13, S18, S28, S30] using LIWC categories like affect, emotions, etc. Other tasks included assessing characteristics, such as goal orientation and sociability [S5, S10, S23, S26] and evaluating the linguistic features of LLM responses to Stack Overflow questions.
On the other hand, 21 studies employed LIWC’s output indirectly, using it to develop other algorithms or models, or to extract insights that necessitated further processing of LIWC’s outputs. The most common method of using LIWC indirectly was to uncover developers’ personalities [S2, S8, S9, S19, S27, S35]. Indeed, prior research has demonstrated multiple methods of inferring text authors’ personalities using LIWC. Pennebaker & King [60], Yarkoni et al. [30], and Goldbeck et al. [61] all have proposed methods of leveraging LIWC for measuring the personality traits of an author based on the widely used Big-Five Factor personality model [62]. However, it should be noted that while the original studies that employ LIWC categories to measure personality traits have received significant attention [60, 30, 61], their focus has predominantly centered on general rather than domain-specific text, such as software engineering-related content.